OpenStudy (anonymous):

A is a non singular matrix of order 3 and IAI = 5. find I adj A I ?

OpenStudy (anonymous):

hey @angela210793 help

OpenStudy (anonymous):

There is one formula for this you will have to use that..

OpenStudy (anonymous):

My equation editor is not working... I am refreshing the page just a sec..

OpenStudy (anonymous):

\[\large \color{green}{{\left| Adj A \right| = \left| A \right|^{n-1}}, \quad n= Order \; Of \; Matrix}\] here n= 3, \[\large \color{blue}{\left| AdjA \right| = 5^2 = 25}\]